# Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands

### v1

@@ II S2: Superfluid Weight Calculation for Singular Two-Band Models
To exhibit a $\Delta\ln(a/\Delta)$ behavior, we assume model 17 lies in the linear regime of some fictitious parameter $\lambda$. To incorporate with Eq. 4, we allow the group velocity of the two bands to be different by adding a $\sigma_{0}$ term, giving the modified Hamiltonian
$h(\textbf{k})=v(k_{x}\hat{\sigma}_{x}+k_{y}\hat{\sigma}_{y})+v^{\prime}k\hat{\sigma}_{0}.$ (20)
The two bands have group velocity $v_{c}=v^{\prime}+v$ and $v_{v}=v^{\prime}-v$, respectively, but the quantum metric is still Eq. 18. Now the model has three energy scales $v_{v}k_{\Lambda}$, $v_{c}k_{\Lambda}$ and $\Delta$, from which we define $\lambda=-v_{v}/v_{c}$ and $\kappa=v_{c}k_{\Lambda}/\Delta$. We assume $v^{\prime}\geqslant 0$, so $|\lambda|\leqslant 1$. At the critical value $\lambda=0$, the lower band becomes flat.
Eq. 9 now gives the total geometric superfluid weight
$D_{s}^{\text{geo}}=\frac{\Delta}{4\pi}\frac{1+\lambda}{1-\lambda}\ln\frac{1+a_{\kappa}}{1+a_{\lambda\kappa}},$ (21)
which contains the crossover function $\chi(\lambda,\kappa)$, while the conventional superfluid weight can be shown not containing $\chi(\lambda,\kappa)$. Functions of this model are listed in Table 2.
Table 2: Functions of the Dirac model 20
@@
In the table both $f_{1}^{\text{geo}}$ and $f_{1}^{\text{conv}}\neq 0$, indicating that for $\lambda\neq 0$, in the regime $\lambda\kappa\gg 1$ both $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}\propto\Delta$. As the valence band becomes flat ($\lambda\rightarrow 0$), $D_{s}^{\text{conv}}$ remains linear while $D_{s}^{\text{geo}}$ has a $\Delta\sim\Delta\ln(v_{c}k_{\Lambda}/2\Delta)$ crossover. Analyzing the slope function $f_{1}^{\text{geo}}$, at $\lambda=1$ it is finite, corresponding to the original Dirac model 17; $\lambda=0$ is the critical point when the valence band becomes flat, so it diverges; $\lambda=-1$ is a “non-geometric limit” when the two bands become degenerate so it vanishes.
@@ III S3: Superfluid Weight Calculation for Singular Three-Band Models
$\hat{\lambda}_{x}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\hat{\lambda}_{y}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\hat{\lambda}_{z}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}.$ (22)
for Eq. 3. A realization of this Hamiltonian for $J=1$ in Lieb lattice is shown in Fig. 6(a). The imaginary hoppings ±$\pm im/4$ between $A$ and $C$ atoms open a band gap $m$ between the flat and valence/conduction band. $\hat{\lambda}_{z}$ term makes the model non-bipartite and Lieb theorem no longer applicable [34], but one can still assume a superconducting exact ground state for small $m$.
Using Fourier transform $c_{\textbf{R}\alpha}=\frac{1}{\sqrt{N}}\sum_{\textbf{k}}e^{i\textbf{k}\cdot(\textbf{R}+\textbf{r}_{\alpha})}c_{\textbf{k}\alpha}$, with $\textbf{R}=i\hat{\textbf{x}}+j\hat{\textbf{y}}$ the lattice vectors, $\textbf{r}_{A}=(\frac{1}{2},0)$, $\textbf{r}_{B}=(0,0)$, $\textbf{r}_{C}=(0,\frac{1}{2})$ the atom positions within a unit cell, the model in Fig. 6(a) gives the Bloch Hamiltonian
$h_{lattice}^{J=1}(\textbf{k})=\begin{pmatrix}0&2t\cos\frac{k_{x}}{2}&-im\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\\ 2t\cos\frac{k_{x}}{2}&0&2t\cos\frac{k_{y}}{2}\\ im\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}&2t\cos\frac{k_{y}}{2}&0\end{pmatrix}.$ (23)
Setting $\zeta_{1}=-t$ and expanding it around the gap closing point $\textbf{k}_{M}=(\pi,\pi)$, one obtains Eq. 3 for $J=1$.
One can show that PHS combined with inversion symmetry leads to an important property of the three-band chiral Hamiltonian Eq. 3, that the inter-band quantum metric between the valence and conduction bands vanishes, $g^{+-}_{\mu\nu}(\textbf{k})=0$ (for definition see Eq. II).
@@
$|u_{\pm}(-\textbf{k})\rangle=V|u_{\pm}(\textbf{k})\rangle,$ (24)
where $|u^{*}\rangle$ denotes the column vector with each component of $|u\rangle$ taking complex conjugate. Property $U$ and $V$ in this equation imply that the connection between valence and conduction band, $\mathcal{A}_{\mu}(\textbf{{k}})\equiv\langle u_{+}(\textbf{k})|\partial_{\mu}u_{-}(\textbf{k})\rangle$ is simultaneously an odd and even function of k, i.e. by PHS
@@
$=$ $\langle\partial_{\mu}u_{+}(\textbf{k})|u_{-}(\textbf{k})\rangle=-\mathcal{A}_{\mu}(\textbf{k}),$ (25)
@@
$=$ $\langle u_{+}(\textbf{k})|\partial_{\mu}u_{-}(\textbf{k})\rangle=\mathcal{A}_{\mu}(\textbf{k}).$ (26)
@@
Inspecting Hamiltonian Eq. 3, we find it satisfies PHS $U=\text{diag}(1,1,1)$ and inversion $V=\text{diag}(-1,1,-1)$ for odd $J$, while $U=\text{diag}(1,-1,1)$ and $V=\text{diag}(1,1,1)$ for even $J$. Therefore $g^{+-}_{\mu\nu}(\textbf{k})=0$ is justified.
These symmetries also lead to: (1) Berry curvature of the valence and conduction band are exactly opposite, while the flat band has no Berry curvature. (2) quantum metric of the three bands $g^{0}_{\mu\nu},g^{\pm}_{\mu\nu}$ are related to the inter-band (band-resolved) quantum metric in a way $g^{0}_{\mu\nu}(\textbf{k})=2g^{\pm}_{\mu\nu}(\textbf{k})=-2g^{0\pm}_{\mu\nu}(\textbf{k})$. The analytical calculation can show $g^{0}_{\mu\nu}$ is precisely four times the quantum metric of the two-band model, Eq. 2.
Using Eq. 9, we find the total geometric superfluid weight is
$D_{s}^{\text{geo}}=4\Delta\sum_{{\bf k}}\bigg{(}1-\frac{\Delta}{E_{+,{\bf k}}}\bigg{)}g^{0}_{xx}({\bf k}),$ (27)
where $E_{+,{\bf k}}=\sqrt{\varepsilon_{+,\textbf{k}}^{2}+\Delta^{2}}$, with $\varepsilon_{+,\textbf{k}}=\sqrt{\zeta_{J}^{2}k^{2J}+m^{2}}$. Compared with Eq. 14, the conventional and geometric superfluid weights are similar to the two-band chiral model. The slight difference is here we define $E_{g}=m$, $W=\sqrt{\zeta_{J}^{2}k_{\Lambda}^{2J}+m^{2}}$ and $\lambda=E_{g}/W$, $\kappa=W/\Delta$. Then, one can check the functions $\mathcal{F}_{i}^{\alpha},f_{i}^{\alpha}$ are also similar to those in Table 1, except for some numerical factors.
@@
For case (ii) of model Eq. 3, the geometry is the same as case (i). Eq. 27 is still valid, but with $\varepsilon_{+,\textbf{k}}=c(m)$ since the valence and conduction bands are flattened. Performing the integrals, one obtains
$D_{s}^{\text{geo}}=\frac{J\Delta}{2\pi}\bigg{(}1-\frac{1}{\sqrt{1+t^{2}}}\bigg{)}(-2\ln\lambda_{m}+1-\lambda_{m}^{2}),$ (28)
@@
We also present the calculation for the Lieb lattice model in Ref. [16] (Fig. 6(b)). The Hamiltonian is
$h(\textbf{k})=2t\begin{pmatrix}0&a_{\textbf{k}}^{*}&0\\ a_{\textbf{k}}&0&b_{\textbf{k}}\\ 0&b_{\textbf{k}}^{*}&0\end{pmatrix}$ (29)
@@
Figure 6: Contrasting two types of Lieb lattice model with a band gap. Black dots denote the three atoms $A,B,C$ grouped as a unit cell. (a) A non-bipartite lattice with a gap $m$ opened by imaginary hoppings between atom $A$ and $C$. The arrows indicate hopping integrals from $C$ to $A$, while the reverse directions are their complex conjugate. (b) A bipartite lattice from Ref. [16], with a gap $2\sqrt{2}t\delta$ opened by unequal hoppings $(1\pm\delta)t$ between $A,B$ or $B,C$. (a) has inversion symmetry, but (b) does not.
@@
$\varepsilon_{\pm,\textbf{k}}\approx\pm t\sqrt{8\delta^{2}+(1-\delta^{2})k^{2}},$ (30)
@@
$g^{+-}_{\mu\nu}(\textbf{k})\approx-\frac{\delta^{2}(2\delta_{\mu\nu}-1)}{4(4\delta^{2}+\frac{1-\delta^{2}}{2}k^{2})^{2}},$ (31)
$g^{0\pm}_{\mu\nu}(\textbf{k})\approx-\frac{1}{8(4\delta^{2}+\frac{1-\delta^{2}}{2}k^{2})^{2}}\begin{pmatrix}4\delta^{2}+(1-\delta^{2})k^{2}-(1-\delta^{4})k_{x}^{2}&4\delta^{2}-(1-\delta^{2})^{2}k_{x}k_{y}\\ 4\delta^{2}-(1-\delta^{2})^{2}k_{x}k_{y}&4\delta^{2}+(1-\delta^{2})k^{2}-(1-\delta^{4})k_{y}^{2}\end{pmatrix},$ (32)
In the limit $\delta\rightarrow 0$ there is relation $g^{+-}_{\mu\nu}(\textbf{k})\propto\delta^{2}(\textbf{k})$ and $g^{0\pm}_{\mu\nu}(\textbf{k})\propto(k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu})/k^{4}$. The continuum limits superfluid weight expression is complicated but follows the general expression Eq. 4.
For this model define $E_{g}=2\sqrt{2}t\delta$, $\varepsilon_{\Lambda}=t\sqrt{(1-\delta^{2})k_{\Lambda}^{2}}$, $W=\sqrt{\varepsilon_{\Lambda}^{2}+E_{g}^{2}}$, then as usual $\lambda=E_{g}/W$, $\kappa=W/\Delta$ and keep in mind that $\delta$ is dimensionless. It turns out to be convenient to keep $\delta$ in the functions, and $\mathcal{F}_{i}^{\alpha}(\lambda,\kappa,\delta)$ and $f_{i}^{\alpha}(\lambda,\kappa,\delta)$ are listed in Table 3. We find the functions for the conventional superfluid weight are the same as the two-band and three-band chiral model (Table 1) if set $J=1/(1-\delta^{2})$, and the crossover behaviors are similar too.
@@
IV S4: Short-range Nature of the $\Delta\sim\Delta\ln(W/2\Delta)$ Crossover
The $\Delta\sim\Delta\ln(W/2\Delta)$ crossover for the geometric superfluid weight has a general short-range property, as explained below.
$D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ as functions of $\Delta$ for the two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands are plotted in Fig. 7 (same as Fig. 1), which are obtained from either numerical calculation or analytical results from Table 1. Here, we fix $\mu=0$, which means for large $\Delta/\varepsilon_{\Lambda}$ values, especially in Fig. 7(b), the electron density will deviate from $\nu=1/2$. Fig. 7 resembles the $D_{s}$ vs $\Delta$ relation in Lieb lattice [37].
Notice that in Fig. 7(a), the $D_{s}^{\text{geo}}$ curve for $E_{g}=0$ (i.e. $D_{s}^{\text{geo}}\sim\Delta\ln(\varepsilon_{\Lambda}/2\Delta)$) crosses the $E_{g}=0.1\varepsilon_{\Lambda}$ curve; it crosses all the $E_{g}\neq 0$ curves, which can be seen in a larger scale in Fig. 7(b) or Fig. 1 in the main text. The crossing point $\Delta_{c}$ is of the order of $E_{g}$, indicating that $D_{s}^{\text{geo}}(E_{g}=0)>D_{s}^{\text{geo}}(E_{g}\neq 0)$ only when $\Delta<E_{g}$.
Figure 7: $D_{s}$ vs $\Delta$ plot for the two-band model (using Eq. 4 and Table 1), for $m=0$ (blue), 0.1 (orange), 0.5 (green). The parameters are $\zeta_{J}=k_{\Lambda}=J=1$ so $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}=2$. (a) A smaller scale $0\leqslant\Delta\leqslant 0.15\varepsilon_{\Lambda}$. As the gap closes, the geometric superfluid weight has a $\Delta\sim\Delta\ln(a/\Delta)$ crossover, while the conventional superfluid weight has a $\Delta^{2}\sim\Delta$ crossover. The arrow indicates a cross point. (b) A larger scale $0\leqslant\Delta\leqslant 2\varepsilon_{\Lambda}$.
This short-range property can also be understood if we fix $\Delta$ and consider the change of $D_{s}^{\text{geo}}$ vs $E_{g}$, as shown in Fig. 4. Then $D_{s}^{\text{geo}}$ saturates immediately as $E_{g}$ approaches $\Delta$. When the band gap completely closes, this saturation is necessary for a finite $D_{s}^{\text{geo}}$. Otherwise, consider the intermediate parameter regime $\Delta\ll E_{g}\ll W$, the $-2\ln\lambda$ piece in $f_{1}^{\text{geo}}(\lambda)$ in Table 1 implies that $D_{s}^{\text{geo}}$ follows the law
$D_{s}^{\text{geo}}\propto\Delta\ln\frac{W}{E_{g}},$ (33)

### v2

@@ II S2: Superfluid Weight Calculation for Singular Two-Band Models
To exhibit a $\Delta\ln(a/\Delta)$ behavior, we assume model (S10) lies in the linear regime of some fictitious parameter $\lambda$. To incorporate with Eq. 4 in MT, we allow the group velocity of the two bands to be different by adding a $\sigma_{0}$ term, giving the modified Hamiltonian
$h(\textbf{k})=v(k_{x}\hat{\sigma}_{x}+k_{y}\hat{\sigma}_{y})+v^{\prime}k\hat{\sigma}_{0}.$ (S13)
The two bands have group velocity $v_{c}=v^{\prime}+v$ and $v_{v}=v^{\prime}-v$, respectively, but the quantum metric is still Eq. (S11). Now the model has three energy scales $v_{v}k_{\Lambda}$, $v_{c}k_{\Lambda}$ and $\Delta$, from which we define $\lambda=-v_{v}/v_{c}$ and $\kappa=v_{c}k_{\Lambda}/\Delta$. We assume $v^{\prime}\geqslant 0$, so $|\lambda|\leqslant 1$. At the critical value $\lambda=0$, the lower band becomes flat.
Eq. (S2) now gives the total geometric superfluid weight
$D_{s}^{\text{geo}}=\frac{\Delta}{4\pi}\frac{1+\lambda}{1-\lambda}\ln\frac{1+a_{\kappa}}{1+a_{\lambda\kappa}},$ (S14)
which contains the cross-over function $\chi(\lambda,\kappa)$, while the conventional superfluid weight can be shown not containing $\chi(\lambda,\kappa)$. Functions of this model are listed in Table 2.
Table 2: Functions of the Dirac model (S13)
@@
In the table both $f_{1}^{\text{geo}}$ and $f_{1}^{\text{conv}}\neq 0$, indicating that for $\lambda\neq 0$, in the regime $\lambda\kappa\gg 1$ both $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}\propto\Delta$. As the valence band becomes flat ($\lambda\rightarrow 0$), $D_{s}^{\text{conv}}$ remains linear while $D_{s}^{\text{geo}}$ has a $\Delta\sim\Delta\ln(v_{c}k_{\Lambda}/2\Delta)$ cross-over. Analyzing the slope function $f_{1}^{\text{geo}}$, at $\lambda=1$ it is finite, corresponding to the original Dirac model (S10); $\lambda=0$ is the critical point when the valence band becomes flat, so it diverges; $\lambda=-1$ is a “non-geometric limit” when the two bands become degenerate so it vanishes.
@@ III S3: Superfluid Weight Calculation for Singular Three-Band Models
$\hat{\lambda}_{x}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\,\,\,\hat{\lambda}_{y}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\,\,\,\hat{\lambda}_{z}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}.$ (S15)
for Eq. 3 in MT. A realization of this Hamiltonian for $J=1$ in Lieb lattice is shown in Fig. S1(a). The imaginary hoppings ±$\pm im/4$ between $A$ and $C$ atoms open a band gap $m$ between the flat and valence/conduction band. $\hat{\lambda}_{z}$ term makes the model non-bipartite and Lieb theorem no longer applicable [39], but one can still assume a superconducting exact ground state for small $m$.
Using Fourier transform $c_{\textbf{R}\alpha}=\frac{1}{\sqrt{N}}\sum_{\textbf{k}}e^{i\textbf{k}\cdot(\textbf{R}+\textbf{r}_{\alpha})}c_{\textbf{k}\alpha}$, with $\textbf{R}=i\hat{\textbf{x}}+j\hat{\textbf{y}}$ the lattice vectors, $\textbf{r}_{A}=(\frac{1}{2},0)$, $\textbf{r}_{B}=(0,0)$, $\textbf{r}_{C}=(0,\frac{1}{2})$ the atom positions within a unit cell, the model in Fig. S1(a) gives the Bloch Hamiltonian
$h_{lattice}^{J=1}(\textbf{k})=\begin{pmatrix}0&2t\cos\frac{k_{x}}{2}&-im\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}\\ 2t\cos\frac{k_{x}}{2}&0&2t\cos\frac{k_{y}}{2}\\ im\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}&2t\cos\frac{k_{y}}{2}&0\end{pmatrix}.$ (S16)
Setting $\zeta_{1}=-t$ and expanding it around the gap closing point $\textbf{k}_{M}=(\pi,\pi)$, one obtains Eq. 3 of MT for $J=1$.
One can show that PHS combined with inversion symmetry leads to an important property of the three-band chiral Hamiltonian Eq. 3, that the inter-band quantum metric between the valence and conduction bands vanishes, $g^{+-}_{\mu\nu}(\textbf{k})=0$ (for definition see Eq. (II)).
@@
$|u_{\pm}(-\textbf{k})\rangle=V|u_{\pm}(\textbf{k})\rangle,$ (S17)
where $|u^{*}\rangle$ denotes the column vector with each component of $|u\rangle$ taking complex conjugate. Properties $U$ and $V$ in this equation imply that the connection between valence and conduction band, $\mathcal{A}_{\mu}(\textbf{{k}})\equiv\langle u_{+}(\textbf{k})|\partial_{\mu}u_{-}(\textbf{k})\rangle$ is simultaneously an odd and even function of k, i.e. by PHS
@@
$=$ $\langle\partial_{\mu}u_{+}(\textbf{k})|u_{-}(\textbf{k})\rangle=-\mathcal{A}_{\mu}(\textbf{k}),$ (S18)
@@
$=$ $\langle u_{+}(\textbf{k})|\partial_{\mu}u_{-}(\textbf{k})\rangle=\mathcal{A}_{\mu}(\textbf{k}).$ (S19)
@@
Inspecting Hamiltonian Eq. 3 in MT, we find it satisfies PHS $U=\text{diag}(1,1,1)$ and inversion $V=\text{diag}(-1,1,-1)$ for odd $J$, while $U=\text{diag}(1,-1,1)$ and $V=\text{diag}(1,1,1)$ for even $J$. Therefore $g^{+-}_{\mu\nu}(\textbf{k})=0$ is justified.
These symmetries also lead to: (1) Berry curvature of the valence and conduction band are exactly opposite, while the flat band has no Berry curvature. (2) quantum metric of the three bands $g^{0}_{\mu\nu},g^{\pm}_{\mu\nu}$ are related to the inter-band (band-resolved) quantum metric in a way $g^{0}_{\mu\nu}(\textbf{k})=2g^{\pm}_{\mu\nu}(\textbf{k})=-2g^{0\pm}_{\mu\nu}(\textbf{k})$. The analytical calculation can show $g^{0}_{\mu\nu}$ is precisely four times the quantum metric of the two-band model, Eq. 2 in MT.
Using Eq. (S2), we find the total geometric superfluid weight is
$D_{s}^{\text{geo}}=4\Delta\sum_{{\bf k}}\bigg{(}1-\frac{\Delta}{E_{+,{\bf k}}}\bigg{)}g^{0}_{xx}({\bf k}),$ (S20)
where $E_{+,{\bf k}}=\sqrt{\varepsilon_{+,\textbf{k}}^{2}+\Delta^{2}}$, with $\varepsilon_{+,\textbf{k}}=\sqrt{\zeta_{J}^{2}k^{2J}+m^{2}}$. Compared with Eq. (S7), the conventional and geometric superfluid weights are similar to the two-band chiral model. The slight difference is here we define $E_{g}=m$, $W=\sqrt{\zeta_{J}^{2}k_{\Lambda}^{2J}+m^{2}}$ and $\lambda=E_{g}/W$, $\kappa=W/\Delta$. Then, one can check the functions $\mathcal{F}_{i}^{\alpha},f_{i}^{\alpha}$ are also similar to those in Table 1 of MT, except for some numerical factors.
@@
For case (ii) of model Eq. 3, the geometry is the same as case (i). Eq. (S20) is still valid, but with $\varepsilon_{+,\textbf{k}}=c(m)$ since the valence and conduction bands are flattened. Performing the integrals, one obtains
$D_{s}^{\text{geo}}=\frac{J\Delta}{2\pi}\bigg{(}1-\frac{1}{\sqrt{1+t^{2}}}\bigg{)}(-2\ln\lambda_{m}+1-\lambda_{m}^{2}),$ (S21)
@@
We also present the calculation for the Lieb lattice model in Ref. [20] (Fig. S1(b)). The Hamiltonian is
$h(\textbf{k})=2t\begin{pmatrix}0&a_{\textbf{k}}^{*}&0\\ a_{\textbf{k}}&0&b_{\textbf{k}}\\ 0&b_{\textbf{k}}^{*}&0\end{pmatrix}$ (S22)
@@
Figure S1: Contrasting two types of Lieb lattice model with a band gap. Black dots denote the three atoms $A,B,C$ grouped as a unit cell. (a) A non-bipartite lattice with a gap $m$ opened by imaginary hoppings between atom $A$ and $C$. The arrows indicate hopping integrals from $C$ to $A$, while the reverse directions are their complex conjugate. (b) A bipartite lattice from Ref. [20], with a gap $2\sqrt{2}t\delta$ opened by unequal hoppings $(1\pm\delta)t$ between $A,B$ or $B,C$. (a) has inversion symmetry, but (b) does not.
@@
$\varepsilon_{\pm,\textbf{k}}\approx\pm t\sqrt{8\delta^{2}+(1-\delta^{2})k^{2}},$ (S23)
@@
$g^{+-}_{\mu\nu}(\textbf{k})\approx-\frac{\delta^{2}(2\delta_{\mu\nu}-1)}{4(4\delta^{2}+\frac{1-\delta^{2}}{2}k^{2})^{2}},$ (S24)
$g^{0\pm}_{\mu\nu}(\textbf{k})\approx-\frac{1}{8(4\delta^{2}+\frac{1-\delta^{2}}{2}k^{2})^{2}}\begin{pmatrix}4\delta^{2}+(1-\delta^{2})k^{2}-(1-\delta^{4})k_{x}^{2}&4\delta^{2}-(1-\delta^{2})^{2}k_{x}k_{y}\\ 4\delta^{2}-(1-\delta^{2})^{2}k_{x}k_{y}&4\delta^{2}+(1-\delta^{2})k^{2}-(1-\delta^{4})k_{y}^{2}\end{pmatrix},$ (S25)
In the limit $\delta\rightarrow 0$ there is relation $g^{+-}_{\mu\nu}(\textbf{k})\propto\delta^{2}(\textbf{k})$ and $g^{0\pm}_{\mu\nu}(\textbf{k})\propto(k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu})/k^{4}$. The continuum limit of the superfluid weight expression is complicated but follows the general expression Eq. 4 in MT.
For this model define $E_{g}=2\sqrt{2}t\delta$, $\varepsilon_{\Lambda}=t\sqrt{(1-\delta^{2})k_{\Lambda}^{2}}$, $W=\sqrt{\varepsilon_{\Lambda}^{2}+E_{g}^{2}}$, then as usual $\lambda=E_{g}/W$, $\kappa=W/\Delta$ and keep in mind that $\delta$ is dimensionless. It turns out to be convenient to keep $\delta$ in the functions, and $\mathcal{F}_{i}^{\alpha}(\lambda,\kappa,\delta)$ and $f_{i}^{\alpha}(\lambda,\kappa,\delta)$ are listed in Table 3. We find the functions for the conventional superfluid weight are similar to those of the two-band chiral model (Table 1 of MT) if set $J=1/(1-\delta^{2})$, and the cross-over behaviors are similar too.
@@
IV S4: Short-range Nature of the $\Delta\sim\Delta\ln(W/2\Delta)$ cross-over
The $\Delta\sim\Delta\ln(W/2\Delta)$ cross-over for the geometric superfluid weight has a general short-range property, as explained below.
$D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ as functions of $\Delta$ for the two-band model Eq. 1 of MT are plotted in Fig. S2 (same as Fig. 1 in MT), which are obtained from either numerical calculation or analytical results from Table 1 in MT. Here, we fix $\mu=0$, which means for large $\Delta/\varepsilon_{\Lambda}$ values, especially in Fig. S2(b), the electron density will deviate from $\nu=1/2$. Fig. S2 resembles the $D_{s}$ vs $\Delta$ relation in Lieb lattice [42].
Notice that in Fig. S2(a), the $D_{s}^{\text{geo}}$ curve for $E_{g}=0$ (i.e. $D_{s}^{\text{geo}}\sim\Delta\ln(\varepsilon_{\Lambda}/2\Delta)$) crosses the $E_{g}=0.1\varepsilon_{\Lambda}$ curve; it crosses all the $E_{g}\neq 0$ curves, which can be seen in a larger scale in Fig. S2(b) or Fig. 1 of MT. The crossing point $\Delta_{c}$ is of the order of $E_{g}$, indicating that $D_{s}^{\text{geo}}(E_{g}=0)>D_{s}^{\text{geo}}(E_{g}\neq 0)$ only when $\Delta<E_{g}$.
Figure S2: $D_{s}$ vs $\Delta$ plot for the two-band model (using Eq. 4 and Table 1 of MT), for $m=0$ (blue), 0.1 (orange), 0.5 (green). The parameters are $\zeta_{J}=k_{\Lambda}=J=1$ so $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}=2$. (a) A smaller scale $0\leqslant\Delta\leqslant 0.15\varepsilon_{\Lambda}$. As the gap closes, the geometric superfluid weight has a $\Delta\sim\Delta\ln(a/\Delta)$ cross-over, while the conventional superfluid weight has a $\Delta^{2}\sim\Delta$ cross-over. The arrow indicates a cross point. (b) A larger scale $0\leqslant\Delta\leqslant 2\varepsilon_{\Lambda}$.
This short-range property can also be understood if we fix $\Delta$ and consider the change of $D_{s}^{\text{geo}}$ vs $E_{g}$, as shown in Fig. 4 of MT. Then $D_{s}^{\text{geo}}$ saturates immediately as $E_{g}$ approaches $\Delta$. When the band gap completely closes, this saturation is necessary for a finite $D_{s}^{\text{geo}}$. Otherwise, consider the intermediate parameter regime $\Delta\ll E_{g}\ll W$, the $-2\ln\lambda$ piece in $f_{1}^{\text{geo}}(\lambda)$ in Table 1 of MT implies that $D_{s}^{\text{geo}}$ follows the law
$D_{s}^{\text{geo}}\propto\Delta\ln\frac{W}{E_{g}},$ (S26)